OpenStudy (anonymous):

rewrite the following equation in standard form 9-7x=(4x-3)^2+5 A) 16x^2-17x+5=0 B) 16X^2-17x-5=0 C) 16x^2-31x-5=0 D) 16x^2-31x+5=0

OpenStudy (misty1212):

\[9-7x=(4x-3)^2+5\]?

OpenStudy (misty1212):

you gotta multiply out on the right first do you know how to do that?

OpenStudy (anonymous):

no :( I honestly have no clue how to handle this one

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

first job it to square \[9-7x=(4x-3)(4x-3)+5\] \[9-7x=16x^2-24x+9+5\] \[9-7x=16x^2-24x+14\]

OpenStudy (misty1212):

then add \(7x\) to both sides and subtract \(9\) from both sides to get it equal to zero

OpenStudy (misty1212):

you got it now?

OpenStudy (anonymous):

im trying to figure it out right now

OpenStudy (misty1212):

k

OpenStudy (anonymous):

its 16x^2-17x+5=0 So B Right @misty1212

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

oh not that is A

OpenStudy (misty1212):

your answer was right but your choice was wrong it is \[16x^2-7x+5\] like you said, but the answer choice is A

OpenStudy (anonymous):

lol I meant A thanks @misty1212

OpenStudy (misty1212):

your welcome

OpenStudy (misty1212):

i mean "you're welcome"